[Ncep.nhc.nco_contacts] An Update for EVeryone: TSB Morning Rounds - Fri Jul 28, 2017
pablo.santos at noaa.gov
Fri Jul 28 13:47:28 UTC 2017
*As a follow up to this morning rounds:*
1) TSB/NCO will be performing the NetApp switch today. This means mount
points in all workstations will have to be changed and workstations
rebooted. Goes without saying operational ones will go first in both HSU
and TAFB. Matt Sardi will be coordinating with you to do so at a time of
Those of you with workstations on your desks outside ops accessing
model/sat data will be last ;) we might or not get to you today. But if
not today, certainly by Monday assuming everything goes smoothly.
Matt and Doug will be on call over the weekend to attend to any issues
that come up as a result of the switch. Yes we debated why not hold back
til Monday. The problem with that is we will lose another 2 days meaning
that before we can switch we would need to spend another day Monday just
syncing up old and new mount points before we can switch. So we decided
to go for it now and commit to do what we have to do to support as
smooth of a transition as we can.
For now, be aware of this and let us know of any issues. We still think
the risk of going with the switch today is minimal.
2) We have been receiving complaints from partners/users such as JTWC,
WeatherPredict (FSSE), and others of not being able to access guidance
data from us or getting data that is only partially complete etc.
We are working these issues closely. The JTWC issue has been resolved.
The others we believe we have identified the source of the problems but
we are still testing our hypothesis on that. Should we confirm it, and
after the NetApp migration is completed, we will address the fixes for
those with a very high priority RFC (immediate) next week.
3) Recurring issues with NWPS in TAFB - The model runs fine in the
workstation we are using and ingests back in AWIPS. The problem appears
to be forecasters do not have a reliable way to tell whether a previous
run is complete sending a new one causing now both the previous and new
one to fail. We are looking at coming up with a means to better monitor
the model run status for the forecasters to be able to better tell when
they can send a new request or not.
That is it for now.
On 7/28/17 8:13 AM, TSB Admin wrote:
> Here is today's summary of NHC computer operations:
> --- PS
> 1.) Issues since accessing ATCF from TAFB yesterday. M. Sardi/Doug
> Gear dealt with it. Related to NetApp migration.
> 2.) After successfully receiving the 12Z EPAC, 18Z ATL, and 00Z EPAC
> yesterday in AWIPS the 06Z ATL never came back. It appears overnight a
> second run was submitted before the prev one completed causing problems.
> We are looking at this now (C. Mello and D. Gaer) to reset things.
> Meanwhile, I have asked C. Mello to look for a better way to give TAFB
> a way to monitor the status of a current run before attempting to
> lunch another. More to follow from Chris on this.
> 1.) One CAVE session on the AWIPS-2 at the JHT desk refused to respond
> to the pan/zoom of the mouse. Another CAVE session on the same
> workstation had no such problem. I will have Chris Mello follow this
> up. We need to stress to the staff the need for them to open a ticket
> on the spot when it happens to given NCF the chance to collect logs
> and troubleshoot these problems.
> 2.) FSSE not coming at all. Monica is already working this problem.
> *Other items:**NHC RFCs:*
> 1.) NetApp Migration continues today with big switch taking place
> today. Will need to monitor carefully over the weekend for any
> potential problems.
Dr. Pablo Santos
Technology and Science Branch
National Hurricane Center, Miami, FL
Twitter: @NWSNHC @NHC_Atlantic @NHC_Pacific @NHC_TAFB @NHC_Surge
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Ncep.nhc.nco_contacts